OpenStudy (anonymous):

what does number does h stand for? 4h+6=30

OpenStudy (anonymous):

4h + 6 = 30 -6 -6 4h = 24 h = 6

OpenStudy (anonymous):

yes thats corrrect 6

OpenStudy (anonymous):

thank you